# CHaRNet: Conditioned Heatmap Regression for Robust Dental Landmark Localization

### v1

@@ 4.3.1 Dataset splitting
4.3.2 Training details
All base models and CHaR versions in the results section were trained by minimizing the MSE loss 3 and the combined loss 5, respectively, using the Adam optimizer [64] for a total of 100 epochs using a batch size of 16. The initial learning rate is set to 0.005 and is progressively reduced using the cosine learning rate decay scheduler. To prevent overfitting, weight decay was set to 0.003, serving as a regularization term. Additionally, dropout rates were tuned between 0.3 and 0.5 across different model configurations to enhance generalization. Pytorch [65] was used as the DL framework to conduct all experiments. With respect to hyper-parameters controlling the relative importance in the final loss value in the CHaR versions, $\lambda_{\text{reg}}$ and $\lambda_{\text{cls}}$ are set to 0.001 and 1 respectively, given the natural bigger quantity of the landmark localization loss.
@@ 5.2 Results
The quantitative 3D landmark detection results for the five point cloud base models, namely PointNet++, DGCNN, PointTransformer, PointNeXt, and PointMLP with and without CHaR, are compared in Table 5 and Table 6, using MEDE and MSR as evaluation metrics, respectively. Three key observations can be made. First, the CHaR-augmented versions consistently outperform their base counterparts, demonstrating the effectiveness of the proposed method. Second, the models perform well on "normal" or "common" dental models (complete sets of teeth), whereas cases involving missing teeth pose a significantly more challenging task for automatic 3D landmark detection. This underscores the importance of evaluating landmark detection systems in scenarios that involve missing teeth. Finally, the CHaR-augmented version of PointMLP (CHaRNet) emerges as the best-performing method in all comparisons.
In terms of MEDE, CHaRNet achieves a remarkable 0.51 mm in normal dental models, 1.28 mm in all dentition models, and 3.94 mm when averaged across all dentition types. For MSR, it achieves an impressive 87.06% in normal dental models, 82.40% in all dentition models, and 68.95% when averaged across all dentition types.
@@ 6 Discussion
The results demonstrate the effectiveness of the proposed CHaR module in improving 3D landmark detection for point cloud-based models, particularly in challenging scenarios involving incomplete dental models. This section discusses the implications of these findings, analyzes performance differences between models, and highlights limitations and future directions.
6.1 Impact of the CHaR module
@@
The consistent improvement of CHaR-augmented models over their base counterparts on both the MEDE and the MSR metrics underscores the effectiveness of the CHaR module. By improving the ability to encode local geometric information, the CHaR module appears to mitigate the challenges posed by missing teeth. This is particularly evident in cases where the baseline models struggle, as highlighted in Figure 8. The integration of the CHaR module enables models to better capture the subtle geometric cues necessary for accurate landmark location in scenarios with incomplete dental models.
6.2 Performance Across Dentition Types
The superior performance of all models in "normal" dental models compared to those with missing teeth emphasizes the inherent difficulty of detecting landmarks in incomplete geometries. CHarNet achieves a remarkable MEDE and MSR of 0.51 mm and 87.06%, respectively, on normal dental models, demonstrating its capability to handle simpler cases with precision. However, the increase in error (3.94 mm MEDE) and the reduced success ratio (68.95% MSR) when macroaveraging with more complex dental models highlight the need for continued advancements to effectively address this challenge. This result reflects the dependency of landmark detection on complete structural patterns, which are disrupted in the absence of teeth.
6.3 Comparison Across Models
Among the models evaluated, PointMLP-CHaR consistently achieves the best performance in both MEDE and MSR, suggesting that the architecture of PointMLP is particularly suited to the problem. Other models, such as PointNet++ and PointTransformer, show significant improvement with the CHaR module, though they lag behind PointMLP-CHaR. This disparity may be attributed to differences in how each architecture processes local and global features within point clouds.
6.4 Qualitative Observations
The qualitative analysis in Figure 8 further illustrates the robustness of the CHaR-augmented models in handling complex cases. The visualizations reveal that, while base models frequently misplace landmarks in regions with missing teeth, CHaR-augmented models demonstrate improved localization accuracy. These findings suggest that the CHaR module enhances the model’s ability to infer plausible landmark positions, even in the absence of direct geometric cues.
6.5 Limitations and Future Work
Despite the promising results, certain limitations warrant discussion. First, while the CHaR module significantly improves performance, the MEDE and MSR values in challenging cases with missing teeth indicate room for further enhancement. Future work could explore additional modules or hybrid approaches to address the difficulties posed by incomplete geometries. Second, reliance on specific architectural features, such as those in PointMLP, suggests that the CHaR module may not generalize equally well across all types of models, indicating the need for further adaptation depending on the model used. Investigating how to adapt the CHaR module to a broader range of architectures is a potential avenue for research. Finally, while the presented results focus on the accuracy of landmark detection, future studies could evaluate the efficiency and scalability of the module for larger point clouds or real-time applications.
@@ 7 Conclusion
This work introduces the CHaRNet and the CHaR module, a novel approach to enhance 3D landmark detection in point cloud-based models, particularly for dental applications. The results demonstrate that the CHaR module significantly improves performance across multiple architectures, as evidenced by consistent reductions in MEDE and increases in MSR for both "normal" dental models and cases involving missing teeth. By effectively grouping landmarks associated with each tooth, the CHaR module enables the model to adapt dynamically based on the tooth’s presence or absence, assigning landmarks to a null point if the tooth is missing or accurately localizing them within the dentition if the tooth is present. This capability addresses a key challenge in dental landmark detection, where structural irregularities and missing teeth often hinder model performance.
The superior performance of the CHaR-augmented version in every model, and particularly in PointMLP highlights its suitability for incorporating the proposed module, achieving state-of-the-art results with a MEDE of 0.51 mm in normal dental models and a notable improvement to 1.28 mm across all dentition types. These findings validate the efficacy of the CHaR module in improving the accuracy and robustness of landmark detection, particularly in challenging scenarios.

### v4

@@ 4.3 Implementation details
Training details
All base models and CHaR versions in the results section were trained by minimizing the MSE loss (Formula 3) and the combined loss (Formula 5), respectively, using the Adam optimizer [62] for a total of 100 epochs using a batch size of 16. The initial learning rate is set to 0.005 and is progressively reduced using the cosine learning rate decay scheduler. To prevent overfitting, weight decay was set to 0.003, serving as a regularization term. Additionally, dropout rates were tuned between 0.3 and 0.5 across different model configurations to enhance generalization. Pytorch [63] was used as the DL framework to conduct all experiments. With respect to hyper-parameters controlling the relative importance in the final loss value in the CHaR versions, $\lambda_{\text{reg}}$ and $\lambda_{\text{cls}}$ are set to 0.001 and 1 respectively, given the natural bigger quantity of the landmark localization loss.
@@ 5.2 Results
The quantitative 3D landmark detection results for the five point cloud base models, namely PointNet++, DGCNN, PointTransformer, PointNeXt, and PointMLP with and without CHaR, are compared in Table 5 and Table 6, using MEDE and MSR as evaluation metrics, respectively. Three key observations can be made. First, the CHaR-augmented versions consistently outperform their base counterparts, demonstrating the effectiveness of the proposed method. Second, the models perform well on typical dental models, whereas cases involving missing teeth pose a significantly more challenging task for automatic 3D landmark detection. This underscores the importance of evaluating landmark detection systems in scenarios that involve missing teeth. Finally, the CHaR-augmented version of PointMLP (CHaRNet) emerges as the best-performing method in all comparisons.
In terms of MEDE, CHaRNet achieves a remarkable 0.51 mm in typical dental models, 1.28 mm in all dentition models, and 3.94 mm when averaged across all dentition types. For MSR, it achieves an impressive 87.06% in typical dental models, 82.40% in all dentition models, and 68.95% when averaged across all dentition types.
@@ 6 Discussion
The results demonstrate the effectiveness of the proposed CHaR module in improving 3D landmark detection for point cloud-based models, particularly in challenging scenarios involving incomplete dental models. Across multiple architectures, CHaR-augmented models consistently outperform their baseline counterparts on both MEDE and MSR, underscoring the impact of enhancing the encoding of local geometric information. This improvement is especially evident in cases with missing teeth, where baseline models often fail to capture the subtle cues necessary for accurate landmark location.
When comparing performance across different dentition types, all models perform better on typical dental models than on incomplete ones. For instance, CHarNet achieves a MEDE of 0.51mm and an MSR of 87.06% on typical models, highlighting its ability to handle simpler cases with precision. However, macroaveraged results that include more complex scenarios exhibit a higher error (3.94mm MEDE) and lower success ratio (68.95% MSR), suggesting that the absence of certain teeth disrupts key structural patterns used for landmark detection. Although the CHaR module significantly narrows this performance gap, there remains a clear need for methods better equipped to handle incomplete geometries.
Among the point cloud learning models evaluated, PointMLP demonstrates the most pronounced gains with the addition of CHaR module, indicating that PointMLP’s architecture may be particularly compatible with CHaR’s approach to capturing and processing local features. Other architectures such as PointNet++ and PointNeXt also benefit from the inclusion of the CHaR module, but do not reach the same level of improvement. This disparity may be attributed to differences in how each model balances local and global feature extraction.
Qualitative analyses reinforce these quantitative findings. Visual inspections, as shown in Figure 8, reveal that while baseline models frequently misplace landmarks in regions with missing teeth, CHaR-augmented models are more robust, even when direct geometric cues are limited. The inferred landmark placements in these complex regions highlight the CHaR module’s capability to reason about plausible geometries in the face of missing data.
Despite these promising results, certain limitations suggest directions for future work. Although the CHaR module significantly enhances performance in incomplete cases, current MEDE and MSR values leave room for improvement. Approaches that combine the CHaR module with additional processing steps or other network modules could be explored to address more severe or irregular forms of dental model incompleteness. Furthermore, because the CHaR module appears to rely on specific architectural features of PointMLP, efforts to generalize it to a broader range of network designs could expand its utility. Finally, while the focus here is on landmark detection accuracy, real-time processing and scalability to larger point clouds remain open questions that merit further investigation.
@@ 7 Conclusion
This work introduces the CHaRNet and the CHaR module, a novel approach to enhance 3D landmark detection in point cloud-based models, particularly for dental applications. The results demonstrate that the CHaR module significantly improves performance across multiple architectures, as evidenced by consistent reductions in MEDE and increases in MSR for both typical dental models and cases involving missing teeth. By effectively grouping landmarks associated with each tooth, the CHaR module enables the model to adapt dynamically based on the tooth’s presence or absence, assigning landmarks to a null point if the tooth is missing or accurately localizing them within the dentition if the tooth is present. This capability addresses a key challenge in dental landmark detection, where structural irregularities and missing teeth often hinder model performance.
The superior performance of the CHaR-augmented version in every model, particularly in PointMLP, highlights its suitability for incorporating the proposed module, achieving state-of-the-art results with a MEDE of 0.51 mm in typical dental models and a robust performance with 1.28 mm across all dentition types. These findings validate the efficacy of the CHaR module in improving the accuracy and robustness of landmark detection, particularly in challenging scenarios.